How many different ways can 4 people: Andy, Betty, Cindy, and Doug, sit in a row at the opera if Andy and Betty must sit together. I think the answer is 12.

Answer by stanbon(75887)

How many different ways can 4 people: Andy, Betty, Cindy, and Doug, sit in a row at the opera if Andy and Betty must sit together. I think the answer is 12.
--------------------
A/B act like one person.
So there are 3! ways to arrange the "3" people
--
A/B can be arrange in 2! = 2 ways.
---
Total arrangements: 3!*2 = 6*2 = 12
----
You are correct.
=====================
Cheers,
Stan H.